□      Exercise: Suppose the demand curve is linear and is given by the equation P = a – bQ where P is price and Q is quantity. What is the consumer surplus if the equilibrium price is P* and equilibrium quantity is Q*?


1
Expert's answer
2021-07-30T04:49:02-0400

Solution


Consumer Surplus =

Total Utility – (Price x Quantity)


Demand is given by : p = a - bQ


Marginal revenue(MR) = "\frac{d(TR)}{dQ}

d(TR)"

"Marginalrevenue(MR)= \n\ndQ"


Where TR="P\times Q=(a−bQ)\times Q"


"=aQ - bQ^2"

Thus, MR = "\frac{d(TR)}{dQ} = a - 2bQ"


Hence, "MR = a - 2bQ"


When Q = 0


"Then,P=a−b\times0=a"


"MR = a - 2b\times0 = a"


Consumer surplus = "aQ−bQ ^ 2 - (a\times 0) aQ−bQ 2−(a×0)"


= "aQ\u2212bQ ^ 2=aQ\u2212bQ^2"

= 0-0 = 0